Status Report

ISS On-Orbit Status 24 Jan 2003

By SpaceRef Editor
January 24, 2003
Filed under , ,
ISS On-Orbit Status 24 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.

More kudos for the crew this morning for yesterday’s flawless job on the SSRMS/robotarm and for the Russian PAO downlink to the closing ceremony of the 31st Russian
National Space Competition in Korolev.

Right after wakeup (1:00am EST) and before breakfast, all crewmembers completed another session with the medical assessment MO-9 of the Russian crew health-monitoring
program. [MO-9 is biochemical urinalysis, conducted regularly every 30 days (and also before and after EVAs), and it is one of five nominal Russian medical tests adopted
by NASA for US crewmembers for PHS (periodic health status) evaluation as part of the “PHS/Without Blood Labs” exam.]

After breakfast, CDR Ken Bowersox and FE-2/SO Don Pettit performed the remainder of the PHS/Without Blood Labs protocol, each taking turns to assist his crewmate as
CMO. [The data from both tests were entered in the IFEP (in-flight examination program) on the MEC (medical equipment computer) by Pettit, and all PHS hardware was
then stowed.]

Bowersox today completed the R.3 software upgrade on EXPRESS Rack 4 (ER4) and its RIC (rack interface controller) laptop, all previous ER upgrades having gone well so
far. This was the third of five such upgrades (ER5 on 1/17; ER1 on 1/20), and there were no impacts on science or continuously powered payloads. [Today’s procedures were
updated since the last load, including a change to correct an improper IP setting command discovered after the ER5 loading (ref. OOS report 1/19).]

FE-1 Nikolai Budarin conducted the third experiment run of the Russian Plasma Crystal-3 (PK-3) payload, activating the turbopump for work chamber evacuation, setting up the
hardware and preparing the video recording, while tagging up with ground specialists. During the run, Nikolai periodically monitored pressure readings and reported result
during comm passes. Finally, the PK-3 hardware was dismantled. [Today’s experiment looked at large particles, searching for mixture modes without voids in the center at
varied discharge power and varied pressures, measuring the residual charge of fine particles after termination of charging (relaxation) at varied pressures, and studying
general wave dispersion and solitary pressure wave dispersion in mixtures, then repeating the first part for comparison.]

In preparation of the second round of the NASA renal (kidney) stone experiment activities by Expedition 6, Don Pettit started his diet log, then unstowed and set up the
experiment hardware. [This long-range preventive medicine investigation features daily random ingestion of either potassium citrate or placebo tablets. It is Dr. Peggy
Whitson’s double-blind research study investigating methods to prevent formation of kidney stones in zero-G. Part of the experiment consists in keeping a metabolic diet log
(food and fluid intake), followed by collection of urine samples several times per day during each week-long session.].

Flight Engineer Budarin performed another periodic in-flight maintenance (IFM) job: removing and replacing FGB battery #3, one of the six 800A storage units in the Zarya
module. The failed 800A was stowed for disposal on Progress. All other FGB batteries are reportedly working nominally.

Don Pettit completed the daily routine maintenance of the SOSh life support system in the SM (incl. ASU toilet system) as well as the daily IMS (inventory management
system) delta file preparation for automated export to MCC-M/TsUP, while Bowersox did the daily Lab payload status checkup (PCG-STES010 & ZCG).

A special teleconference was conducted by the crew with TsUP IMS specialists to discuss disposition of various pieces of equipment for updating the IMS database. [Items
covered included the dismantled Vozdukh vacuum pumps, cables, SM and DC dust filters, video tapes, etc.]

Today also marked the begin of pre-packing activities for ULF-1, kicked off with a teleconference between crew and MCC-H at 6:15am EST. A preliminary pre-pack list has
been uplinked, which highlights onboard items that can already be packed. [The ULF-1 return manifest is already full. However, return candidates can still be identified since
items may drop off the list or priorities may change for return items.]

Bowersox and Pettit also conducted a teleconference from the Lab on EVA tools stowage for ULF-1, going by a detailed Airlock (A/L) equipment layout plan uplinked last night.

At 9:40am, the crew was interviewed live on TV by stations KRQE-TV in Albuquerque, NM, and WTSP-TV in St. Petersburg, FL.

All crewmembers performed their daily physical exercise on TVIS, RED, and VELO with load trainer.

All SSRMS/robotarm objectives yesterday were completed nominally by Sox and Don. Canadarm2 is currently in the S1 radiator viewing position for Monday’s (1/27) survey
of the clearance issue, with MBS (mobile base system) and SSRMS both on KA (keep-alive) power.

With this morning’s final pressurization of the ISS interior with fresh oxygen (O2), all propellant and gaseous O2 transfers from Progress M1-258 (9P) have been completed.
Begin of trash loading of Progress is scheduled 1/27 (Monday), and undocking is still set for 2/1. [Today’s O2 repress showed two minor anomalies, which are currently
being look into by MCC-H: (a) rate of pressure increase was higher than expected; (b) equal amounts of O2 and N2 (nitrogen) were received.]

After Sox’s last FOOT (foot/ground reaction forces during space flight) run on 1/22, POC (Payload Operations Center) reports that this completed three of five scheduled runs.

Integrated testing of the new MSG (microgravity science glovebox) PDC (power distribution controller) at MSFC was successfully completed. Both the PDC and the ESEM3
(exchangeable standard electronic module 3) have arrived at RSC-Energia in Moscow today. RSC-E has delayed the departure of its last airplane to Baikonur/Kazakhstan by
one day to tomorrow (1/25) so that the MSG parts can arrive at the Soyuz rocket launch pad in time for last-minute loading aboard Progress 10P.

The water venting from the WRM (water recovery and management) condensate tank in the Lab and two CWCs (contingency water containers) on 1/22 went very well.
[During the process, CWC #1022 was found to be leaking and will not be used anymore. Only a tiny amount of water escaped harmlessly into the cabin.]

As stipulated by the current “bottleneck” of the Russian SRV-K water processing system with regard to U.S. condensate (due to limited capacity of the CFU condensate feed
unit of the Elektron O2 generator), which requires overboard dumping of excess US water, new work procedures for joint ops between the RS SRV-K and the USOS Lab
CCAA (common cabin air assembly) air conditioner are being discussed by MCC-M and MCC-M.

After the checkout of the secondary sealing surface of a “Gamah” fitting in the line from the recently (1/18) removed and replaced O2 regulator in the A/L to the PCA (pressure
control assembly), its high-pressure side will now be tested during the weekend ahead. It appears to be sealing correctly.

Work has begun today on the replacement of the ECOMM (early communications) antenna mounting plate in the Node starboard hatch (leading to the A/L) with a two-pane
window. [Today’s task for this job, which is on the “job jar” task list, is a leak check of the Node (16 pressure readings over the next 24 hrs) to determine whether the
currently employed VAJ (vacuum access jumper) setup will support small leak rate measurements required to support the leak checks after the window change-out. The
actual removal/installation work, on MCC-H go-ahead, will then take place over this weekend, followed by the hatch window leak check (five pressure readings over 24 hrs),
to verify that the newly installed window panes can support future EVA activities in the A/L.]

Loop-back testing of the DAIU (docked audio interface unit), which had exhibited spurious anomalous behavior during the last Shuttle docking, showed the DAIU itself to be
healthy. Both the ISS side and the Shuttle side of the system have now been exonerated, leaving only the ISS/Shuttle interface as suspect.

As discovered last week (ref. OOS 1/18), both loop B heaters of the NTA (nitrogen tank assembly) of the EATCS (external active thermal control system) on the P1 truss have
unexpectedly failed. Root cause is still unknown. This also lost heater redundancy for the NTA system, which will first be used on Mission 12A.1. Careful heat monitoring is
being employed to keep temperatures above the minimum hardware limit. [Current temperature is fairly stable at -42.5 degC. Should it fall below -48.3 degC, flight attitude
would be changed to LVLH (local vertical/local horizontal) to protect the thermal limits of the hardware.]

MCC-H has determined that the CBCS (centerline berthing camera system) power supply which failed about two weeks ago, causing the UOP3 RPC (utility outlet panel #3
remote power controller) to trip, was indeed the primary supply. [The backup unit will now become the prime unit. A task has been added to the crew’s task list to label and
relocate the newly designated primary power supply.]

Today’s targets for the CEO (crew earth observations program), limited because daylight passes are currently mainly over the oceanic hemisphere, were Sydney, Australia
(center of the city a touch right; ESC [electronic still camera]. Noting the airport runways jutting into Sydney Harbour, and trying to get the entire city in one or two frames),
Perth, Australia (nadir pass over the large Perth-Fremantle port region will require views to both left and right),and Cape Town, South Africa (pass over the Cape
Peninsula: looking left for the large Cape Town urban region. Urban margins are expanding inland fast with a new pulse of urbanization).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 2:35pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On (16 Amp mode). Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is Off. TCCS (trace contaminant control subsystem) is off.
  • MCA (major constituents analyzer) is operating. BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On,
  • SKV-2 is Off.

  • SM Working Compartment: Pressure (mmHg) — 753; temperature (deg C) — 25.8; ppO2 (mmHg) — 169.3; ppCO2 (mmHg) — 2.4 (suspect).
  • SM Transfer Compartment: Pressure (mmHg) — 751; temperature (deg C) — 20.0.FGB Cabin: Pressure (mmHg) — 752; temperature (deg C) — 23.7.
  • Node: Pressure (mmHg) — 747.07; temperature (deg C) — 23.6 (shell); ppO2 (mmHg) — 169.4; ppCO2 (mmHg) — 4.7.
  • U.S. Lab: Pressure (mmHg) — 749.26; temperature (deg C) — 22.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 742.26, temperature (deg C) — 21.9; shell heater temp (deg C) — 22.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 24.7
  • PMA-2: Shell heater temp (deg C) — 14.2

(n/a = data not available)

Propulsion System (PS):

  • Total propellant load available (SM + FGB + Progress) — 3669 kg (8089 lb) [as of 1/23/03]. (Capability: SM — 870 kg; FGB — 6160 kg).

Electrical Power Systems (EPS) — as of 3:00am:

  • Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and BGA 4B both in Autotrack (sun-following) mode.
  • SM batteries (as of 3am): Battery #7 is off line (failed); battery #4 is off line (ROM/capacity restoration mode); all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries (3am): Battery #3 is offline (failed, but was replaced later today); battery #6 is in “Cycle mode”; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup (new patches loaded on both).
  • EXT-1 is On (primary), EXT-2 is off.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-2 MDM is On (primary); PL-1 MDM is off (diagnostic
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational. Lane 1 is down (as of 11/14).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector — U.S. SIGI-1
  • Attitude — U.S. SIGI-1
  • Angular rates — U.S. RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: -179.5 deg, pitch: -8.8 deg., roll: 0 deg]), CMG/Thruster Assist Momentum Management).
  • Solar Beta Angle: 30.8 deg (magnitude decreasing).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operating.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF2 with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.
  • MT: latched at WS4, with KA power.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning,6:26am EST [= epoch]):

  • Mean altitude — 388.6 km
  • Apogee — 392.4 km
  • Perigee — 384.8 km
  • Period — 92.33 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.000561
  • Orbits per 24-hr. day — 15.60
  • Altitude loss — 170 m (mean) in last 24 hours
  • Solar Beta Angle — 30.8 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 23858

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.